OpenStudy (anonymous):

Find a quadratic equation with roots –1 + 4i and –1 – 4i.

OpenStudy (anonymous):

(x+1)^2=-16 x2+2x+17=0

OpenStudy (anonymous):

put it like this: (x-(-1+4i))(x-(-1-4i)=0 (x+1-4i)(x+1+4i)=0 x^2+2x+17=0

OpenStudy (anonymous):

i give you a perfect solution but you dont mention it (x-a)^2=-b2